SUBSTITUTION METHOD 3X - 3Y = 2 EQUATION 1
X + 2 = 3Y EQUATION 2

Answer by venugopalramana(3286)

SUBSTITUTION METHOD 3X - 3Y = 2 EQUATION 1
X + 2 = 3Y EQUATION 2
X=3Y-2
PUT IN EQN.I
3(3Y-2)-3Y=2
9Y-6-3Y=2
6Y=6+2=8
Y=8/6=4/3
X=3*(4/3)-2=4-2=2